OpenStudy (anonymous):

i can't figure out how to model the function rule y= -2x-3 with a table of values

OpenStudy (texaschic101):

y = -2x - 3 now just sub a number in for x and solve for y. Lets make x = 1 y = -2(1) - 3 y = -2 - 3 y = -5 so when x = 1, y = -5 now lets make x = 2 y = -2(2) - 3 y = -4 - 3 y = -7 so when x = 2, y = -7 and you can continue subbing in any number for x to find y....and there you have your table

OpenStudy (aripotta):

this is what your table is gonna look like. do what texaschic said to fill it up :) 

OpenStudy (aripotta):